OpenStudy (anonymous):

PLEASE math final last question will fan and medal! 40. you save 15,000 you place 1/3 in a savings account earning 4.6 APR compounded annually. you then invest one quarter of the remaining balance in a 3 year u.s treasury bond earning a 5.2 APR compounded annually. and the rest in a stock plan. the stock plan increases by 3% in the first year then decreases by 8% the second year and increases 6% the third year what are the balances of each account by the third year and the total gain from the original saved amount (6 points)

OpenStudy (anonymous):

for the savings account i will be having 5722.22$ at the end of third year in the treasury fund 2910.6$ in the stock 7030$

OpenStudy (anonymous):

7030$

OpenStudy (anonymous):

medal plzz

OpenStudy (anonymous):

Sorry its a question i get right if the steps are right too

OpenStudy (anonymous):

can you just put the steps it took to get you the answer??

OpenStudy (anonymous):

@joe.hardy111 @shehryar then ill medal and fan

OpenStudy (anonymous):

\[5000\times (1.046)^3\] for the first one

OpenStudy (anonymous):

\[2000\times (1.052)^3\] for the second

OpenStudy (anonymous):

\[8000\times (1.03)\times (.92)\times (1.06)\]for the third

OpenStudy (anonymous):

u r wrong @satellite73 because the quarter of 10000 is 2500 not 2000

OpenStudy (anonymous):

Ok well sattelite is basically correct but had a small error So there are 3 diff assets within this portfolio. We need to first determine how much each asset is worth in 3 years and additionally we also need to determine the total interest accrued #1 so a 1/3 of the 15,000 goes into a savings acct. \[\frac{15000}{3}=5000\] So 5000 goes into a savings account and earns 4.6% interest for 3 years \[5000*(1+.46)^3=?\] #2 Now whatever is left over 15,000-5000=10,000 So a 1/4 of what is left over gets invested into a treasury bond which earns 5.2% \[\frac{10000}{4}=2500\] \[2500(1+.052)^3=?\] So the treasury bond is worth a total of _____ after 3 years #3 Now the remaining gets invested in stocks The remaining is 10000-2500=7500 We are told that in the first year there is an increase of 13% =(1+.13)=1.13 The second year there is a decrease of 8% =(1-.08)=.92 And the third year it increases by 6% = (1+.06)=1.06 So the amount our stocks are worth in 3 years: \[7500*(1.13)*(.92)*(1.06)=?\]

OpenStudy (anonymous):

Now to determine the total amount of interest earned: We add up the total value of our assets in 3 years time and subtract 15,000 and that will give you the amount earned